# Chaotic discretization theorems for forced linear and nonlinear coupled oscillators

### v1

@@ Abstract
We prove the holding of chaos in the sense of Li-Yorke for a family of four-dimensional discrete dynamical systems that are naturally associated to ODEs systems describing coupled oscillators subject to an external non-conservative force, also giving an example of a discrete map that is Li-Yorke chaotic but not topologically transitive. Analytical results are generalized to a modular definition of the problem and to a system of nonlinear oscillators described by polynomial potentials in one coordinate. We perform numerical simulations looking for a strange attractor of the system; furthermore, we present the bifurcation diagram and perform a bifurcation analysis of the system.
@@ 1 Introduction
The pioneering work Li by Li and Yorke stated for the first time a formal definition of chaos for discrete dynamical systems; their main result, well summed up in the statement “period three implies chaos”, gave a very simple sufficient condition in order to establish the arising of chaotic behavior for one-dimensional maps. Since the formulation of the Li-Yorke theorem, mathematicians put their efforts in order to formulate alternative definitions of discrete chaos, by exploring analytical, metric and topological aspects. In addition to the Li-Yorke chaos, it is worth to cite the Devaney chaos Devaney, Banks, the Kolmogorov-Sinai entropy Kolmogorov1, Sinai, the Block-Coppel chaos Block, the distributional chaos Schweizer. Some of these definitions may show connections one with the other, depending on the topological structure of the space; for example, in Aulbach it is shown that, for continuous mappings of a compact interval into itself, Devaney chaos and Block-Coppel chaos are equivalent and imply Li-Yorke chaos, but not vice versa. This last definition of discrete chaos has been given originally for one-dimensional maps; however, it is easily formulated in higher dimensions.
It is worth to mention some recent results. In Bernardes Li-Yorke chaos is analytically studied for linear operators on Banach and Fréchet spaces, while Wang deeply focus on a stronger definition of dense uniform Li-Yorke chaos. In Zhang1 it is shown how Li-Yorke chaos depends on the topology defining the dynamical system. The topic has been explored also for PDEs Zhang2. See Demirovic, Wei for explicit derivations of Li-Yorke chaos for specific models. An interesting emerging field concerns chaotification techniques for the so-called “anti-control of chaos”; see Moysis for an exhaustive review.
@@
The paper is organized as follows. In Section 2 we recall some basic definitions together with the Li-Yorke theorem and the Marotto-Li-Chen theorem Marotto, Li-Chen1; then, we define the discrete map subject of this work together with a simple notion of discretization. In Section 3 we prove that the system is not topologically transitive for a symmetric choice on the parameters. In Section 4 we prove the emergence of Li-Yorke chaos for proper choices on the parameters and generalize results for a modular definition of the system, leading to three chaotic discretization theorems for forced linear and nonlinear coupled oscillators. In Section 4.1 we comment a possible connection between Li-Yorke chaos and the KAM theory Kolmogorov2, Arnold2, Moser for quasi-integrable systems. In Section 5 we present some numerical results. In Section 5.1 we perform qualitative studies of the orbit of the system, by comparing regular orbits for a symmetric choice on the parameters and the rising of a strange attractor in the general case; in Section 5.2 we present the bifurcation diagram of the system and confirm the rising of period doubling bifurcations through a bifurcation analysis using the MATLAB tool MatContM (https://matcont.sourceforge.io/). Finally, in Section 6 we state our conclusions and propose further developments of this work.
@@ 2 Mathematical preliminaries
We begin this section by recalling the definition of chaos in the sense of Li-Yorke Li, opportunely generalized for higher dimension dynamical systems. In the following we consider a discrete map $f\colon(X,d)\to(X,d)$ defined on a metric space $(X,d)$.
@@
The following fundamental result is an immediate consequence of the main theorem presented in Li.
@@
Li-Yorke theorem holds for one-dimensional continuous map from a real interval into itself, while in general a $3$-periodic point does not imply Li-Yorke chaos for $n$-dimensional maps. A sufficient condition for the onset of Li-Yorke chaos for $n$-dimensional discrete maps has been given in Marotto through the existence of a snap-back repeller fixed point. Marotto theorem presented in Marotto originally exhibited some problems in the proof; it has been well formulated and proved by Li-Chen in Li-Chen1 (see also Li-Chen2 for further discussions about the theorem).
@@
Notice that the discretization notion we define is not a standard discretization algorithm for ODEs systems, such as explicit or implicit Euler discretization (see, e.g., Kazakov); however, our aim is not to obtain a discrete approximation of the ODEs system, but rather to define an alternative discrete version associated to the continuous system and to explore connections between regular and chaotic behavior of both systems. Furthermore, this discrete system can share some properties with the continuous one (e.g. fixed points); on the other hand, it can exhibit a richer dynamics.
Goal of this paper is to show that a large class of continuous dynamical systems has a fixed point discretization that is Li-Yorke chaotic for non trivial choices on the parameters, regardless the integrability or chaoticity of the continuous system. The starting point will be the discrete map defined in Disca1, Disca2; in this work we relax the original constraints on the coefficients and define the Ziegler discrete map as
@@
The map (1) is obtained by discretizing in the above sense the equations of motion of the Ziegler pendulum in the sense of Polekhin Polekhin for a choice on the parameters leading to non-Hamiltonian integrability. We also define the modular Ziegler discrete map as
@@
This last formulation of the problem immediately returns the original physical meaning of the angular coordinates $x$, $z$.
@@ 3 Topological transitivity for the Ziegler map
We recall the definition of chaos in the sense of Devaney Devaney, that is formulated through topological transitivity and density of periodic points.
@@
It is worth to notice that the original definition of Devaney was also considering sensitive dependence on initial data, a typical signature of chaotic systems; this condition has been proved to be redundant in Banks.
In this section we reintroduce the Ziegler map as defined in Disca1, Disca2, that is
@@
In Disca2 it is proved that the map (3) is not Devaney chaotic for $\alpha<0$, $\tilde{\alpha}>0$, $\beta>0$, $\gamma\neq 0$, since it has not dense sets of periodic points; in particular, the map has not even periodic points and the sets of odd periodic points are not dense in $\mathbb{R}^{4}$. We show that the remaining necessary condition for the Devaney chaos is not satisfied for a symmetric choice on the parameters.
@@ 4 Chaotic discretization theorems
and derive our results for the map (25) with arbitrary small values of $\varepsilon_{1}$, $\varepsilon_{2}$; naturally, (25) reduces to (1) for $\varepsilon_{1},\varepsilon_{2}\to 0^{+}$.
We divide the proof in a few steps, by analyzing separately the sufficient conditions stated in Theorem 2.
@@
Fixed points of (1) are known from Disca1; by considering the new definitions given in (1) and (2), we can write them through the solutions of
@@
Let $J$ be the Jacobian of (25). $\forall\varepsilon_{1},\varepsilon_{2}>0$, there exist $\overline{\sigma}(\varepsilon_{1},\varepsilon_{2})\in\mathbb{R}$ and a fixed point $p_{0}$ such that the matrix $J^{T}J(p_{0})$ has all eigenvalues greater than $1$, provided that $|\beta|>\frac{1}{\sqrt{2}}$ and $\sigma>|\overline{\sigma}|$.
@@
Therefore, if $|\beta|>\frac{1}{\sqrt{2}}$, we have $\lambda_{1},\lambda_{\beta},\lambda_{\pm}>1$ $\forall\varepsilon_{1},\varepsilon_{2}>0$ $\forall\sigma>|\overline{\sigma}|$, for a threshold parameter $\overline{\sigma}$ that depends on $\varepsilon_{1},\varepsilon_{2}$. ∎
@@
Given Lemma 3, we can avoid the need to construct a proper neighborhood of $p_{0}$. Furthermore, the map (1) has not even periodic points Disca2, so we can show the existence of a snap-back for the point $p_{0}$ by proving the existence of solutions for the equation $f^{(2)}(q)=p$; since it cannot exists a periodic point of period 2, it follows that $q_{0}\neq p_{0}$.
@@
Then, for $\varepsilon_{1,2}$ sufficiently small, $\exists\overline{\sigma}(\varepsilon_{1},\varepsilon_{2})\in\mathbb{R}$ such that (38) is chaotic in the sense of Li-Yorke $\forall\sigma>|\overline{\sigma}|$.
@@
where $\varepsilon_{1,2}>0$, $\alpha>0$, $\beta>0$, $\gamma\in(-2\pi,0)$, $\sigma>1$ and
@@
Then, for $\varepsilon_{1,2}$ sufficiently small, $\exists\overline{\sigma}(\varepsilon_{1},\varepsilon_{2})\in\mathbb{R}$ such that (40) is chaotic in the sense of Li-Yorke $\forall\sigma>|\overline{\sigma}|$.
@@
and $B^{(j)}$ has at least two distinct real roots. Then, for $\varepsilon_{1,2}$ sufficiently small, $\exists\overline{\alpha}(\varepsilon_{1},\varepsilon_{2})\in\mathbb{R}$ such that (42) is chaotic in the sense of Li-Yorke $\forall\alpha>|\overline{\alpha}|$.
@@
Theorem 5 is particularly relevant in the context of nonlinear oscillators Kovacic, when they are perturbed by periodic external forces, defined by polynomial potentials and are subject to linear velocity-dependent perturbations.
4.1 A comment on scrambled sets and perturbations
@@
In answering this question, a curious connection with the KAM theorem Kolmogorov2, Arnold2, Moser can be conjectured. From KAM theory, we know that a quasi-integrable system exhibits invariant tori in the phase space that are not destroyed by the presence of a perturbation; when acting on the system, the perturbation destroys some tori, while other ones survive but are deformed, according to specific non-resonance conditions. In the context of chaos theory, it might be argued that a perturbation acts on the scrambled set of a Li-Yorke chaotic map in a similar way, in the sense that they could exist specific sufficient conditions such that the scrambled set of a Li-Yorke map survives under a small perturbation. If this is true, it would be possible to deduce chaos for a discrete system in general through perturbative arguments.
@@ 5.1 Looking for a strange attractor
In Figure 3 - 4 we set $\tilde{\alpha}\neq\alpha$. It is recognized the arising of a strange attractor for the system, that exhibits a fractal structure analogous to the Hénon attractor Henon. This attractor is confined to the projection planes $(x,z)$ and $(y,\omega)$, while it is replaced by completely random motion by taking projections on the planes $(x,y)$ and $(z,\omega)$, as shown in Figure 5.
Figure 3: Strange attractor for the projection onto the plane $(x,z)$, non-symmetric case $\tilde{\alpha}\neq-\alpha$. The simulation runs for $10^{6}$ iterations. Parameters: $\alpha=-2.5$, $\tilde{\alpha}=2$, $\beta=0.8$, $\gamma=10$. Initial conditions: $x_{0}=0.2$, $y_{0}=0.1$, $z_{0}=0.8$, $\omega_{0}=0.5$. a) Full attractor. b) Magnification of a). c) Magnification of b).
Figure 4: Strange attractor for the projection onto the plane $(y,\omega)$, non-symmetric case $\tilde{\alpha}\neq-\alpha$. The simulation runs for $10^{6}$ iterations. Parameters: $\alpha=-2.5$, $\tilde{\alpha}=2$, $\beta=0.8$, $\gamma=10$. Initial conditions: $x_{0}=0.2$, $y_{0}=0.1$, $z_{0}=0.8$, $\omega_{0}=0.5$. a) Full attractor. b) Magnification of a). c) Magnification of b).
@@
Now, we briefly focus on the case $\gamma=0$. By recalling the physical origin of the system (1), the term $\gamma$ represents an external non-conservative force acting on an Hamiltonian system; by setting $\gamma=0$, the map comes out from an integrable continuous system whose phase space is foliated in invariant tori, due to Liouville-Arnold theorem Liouville, Arnold1. Then, it is worth to see whether the strange attractor shown in Figure 3 - 4 survives for this unperturbed case.
@@
Figure 6: Strange attractor for the projection onto the plane $(x,z)$, non-symmetric case $\tilde{\alpha}\neq-\alpha$. The simulation runs for $10^{6}$ iterations. Parameters: $\alpha=-2.5$, $\tilde{\alpha}=2$, $\beta=0.8$, $\gamma=0$. Initial conditions: $x_{0}=0.2$, $y_{0}=0.1$, $z_{0}=0.8$, $\omega_{0}=0.5$. a) Full attractor. b) Magnification of a).
Figure 7: Strange attractor for the projection onto the plane $(y,\omega)$, non-symmetric case $\tilde{\alpha}\neq-\alpha$. The simulation runs for $10^{6}$ iterations. Parameters: $\alpha=-2.5$, $\tilde{\alpha}=2$, $\beta=0.8$, $\gamma=0$. Initial conditions: $x_{0}=0.2$, $y_{0}=0.1$, $z_{0}=0.8$, $\omega_{0}=0.5$. a) Full attractor. b) Magnification of a).
In Figure 8 they are shown projections on the planes $(x,z)$ and $(y,\omega)$ of two trajectories with very close initial conditions, giving a further confirmation of the sensitive dependence on initial data for the system.
Figure 8: Strange attractor for two very close initial conditions. Parameters: $\alpha=-3.5$, $\tilde{\alpha}=2$, $\beta=0.8$, $\gamma=-4$. Initial conditions, blue orbit: $x_{0}=1.4$, $y_{0}=0.8$, $z_{0}=2.7$, $\omega_{0}=2.34$. Initial conditions, green orbit: $x_{0}=1.4001$, $y_{0}=0.7999$, $z_{0}=2.7001$, $\omega_{0}=2.3399$. The simulation runs for $10^{5}$ iterations. a) Projection onto the plane $(x,z)$. b) Projection onto the plane $(y,\omega)$, magnification on the center.
5.2 Bifurcation analysis
In this section we perform a bifurcation analysis of the system (2) for $\tilde{\alpha}=-\alpha$. We implement the classical algorithm Strogatz for the construction of the bifurcation diagram. The basic idea of the algorithm is to compute the trajectory of the system for every value of the bifurcation parameter up to a certain iteration step ($n_{it}$), discarding the first iterations ($n_{tran}$) so that the system has actually reached its attractor and any transitory effects are not shown. In principle, we should compute orbits for the four-dimensional system (2); however, for the symmetric case $\tilde{\alpha}=-\alpha$ we can use the following simple result.

### v2

@@ Abstract
We prove the holding of chaos in the sense of Li-Yorke for a family of four-dimensional discrete dynamical systems that are naturally associated to ODEs systems describing coupled oscillators subject to an external non-conservative force, also giving an example of a discrete map that is Li-Yorke chaotic but not topologically transitive. Analytical results are generalized to a modular definition of the problem and to a system of nonlinear oscillators described by polynomial potentials in one coordinate. We perform numerical simulations looking for a strange attractor of the system; furthermore, we perform a bifurcation analysis of the system presenting 1D and 2D bifurcation diagrams, together with spectra of Lyapunov exponents and basins of attraction.
@@ 1 Introduction
The pioneering work Li and Yorke [1975] by Li and Yorke stated for the first time a formal definition of chaos for discrete dynamical systems; their main result, well summed up in the statement “period three implies chaos”, gave a very simple sufficient condition in order to establish the arising of chaotic behavior for one-dimensional maps. Since the formulation of the Li-Yorke theorem, mathematicians put their efforts in order to formulate alternative definitions of discrete chaos, by exploring analytical, metric and topological aspects. In addition to the Li-Yorke chaos, it is worth to cite the Devaney chaos Devaney [2021], Banks et al. [1992], the Kolmogorov-Sinai entropy Kolmogorov [1958], Sinai [1959], the Block-Coppel chaos Block and Coppel [1992], the distributional chaos Schweizer and Smítal [1994]. Some of these definitions may show connections one with the other, depending on the topological structure of the space; for example, in Aulbach and Kieninger [2001] it is shown that, for continuous mappings of a compact interval into itself, Devaney chaos and Block-Coppel chaos are equivalent and imply Li-Yorke chaos, but not vice versa. This last definition of discrete chaos has been given originally for one-dimensional maps; however, it is easily formulated in higher dimensions.
It is worth to mention some recent results. In Bernardes Jr et al. [2015] Li-Yorke chaos is analytically studied for linear operators on Banach and Fréchet spaces, while Li and Wang [2025] deeply focus on a stronger definition of dense uniform Li-Yorke chaos. In Zhang et al. [2023] it is shown how Li-Yorke chaos depends on the topology defining the dynamical system. The topic has been explored also for PDEs Zhang et al. [2025]. See Garić-Demirović et al. [2024], Wei and Zhou [2019] for explicit derivations of Li-Yorke chaos for specific models. An interesting emerging field concerns chaotification techniques for the so-called “anti-control of chaos”; see Moysis et al. [2025] for an exhaustive review.
@@
The paper is organized as follows. In Section 2 we recall some basic definitions together with the Li-Yorke theorem and the Marotto-Li-Chen theorem Marotto [1978], Li and Chen [2003a]; then, we define the discrete map subject of this work together with a simple notion of discretization. In Section 3 we prove that the system is not topologically transitive for a symmetric choice on the parameters. In Section 4 we prove the emergence of Li-Yorke chaos for proper choices on the parameters and generalize results for a modular definition of the system, leading to three chaotic discretization theorems for forced linear and nonlinear coupled oscillators. In Section 4.1 we comment a possible connection between Li-Yorke chaos and the KAM theory Kolmogorov [1979], Arnold [1963], Moser [1966] for quasi-integrable systems. In Section 5 we present some numerical results. In Section 5.1 we perform qualitative studies of the orbit of the system, by comparing regular orbits for a symmetric choice on the parameters and the rising of a confined strange attractor in the general case. In Section 5.2 we present 1D and 2D bifurcation diagrams of the system and confirm the rising of period doubling bifurcations through a bifurcation analysis using the MATLAB tool MatContM (https://matcont.sourceforge.io/). Furthermore, we present spectra of Lyapunov exponents, that confirm its sensitive dependence on initial data for several range of the parameters, together with basins of attraction of the system. Finally, in Section 6 we state our conclusions and propose further developments of this work.
@@ 2 Mathematical preliminaries
We begin this section by recalling the definition of chaos in the sense of Li-Yorke Li and Yorke [1975], opportunely generalized for higher dimension dynamical systems. In the following we consider a discrete map $f\colon(X,d)\to(X,d)$ defined on a metric space $(X,d)$.
@@
The following fundamental result is an immediate consequence of the main theorem presented in Li and Yorke [1975].
@@
Li-Yorke theorem holds for one-dimensional continuous map from a real interval into itself, while in general a $3$-periodic point does not imply Li-Yorke chaos for $n$-dimensional maps. A sufficient condition for the onset of Li-Yorke chaos for $n$-dimensional discrete maps has been given in Marotto [1978] through the existence of a snap-back repeller fixed point. Marotto theorem presented in Marotto [1978] originally exhibited some problems in the proof; it has been well formulated and proved by Li-Chen in Li and Chen [2003a] (see also Li and Chen [2003b] for further discussions about the theorem).
@@
In the vicinity of a snap-back repeller, there exists a set of initial conditions such that they exactly back onto the fixed point, after a certain amount of iterations. Geometrically, this behavior induces to a continuous stretching and folding of the phase space around the snap-back repeller, leading to emergence of chaos.
@@
Notice that the discretization notion we define is not a standard discretization algorithm for ODEs systems, such as explicit or implicit Euler discretization (see, e.g., Kazakov et al. [2024]), and in general it is not appropriate for the numerical resolution of the associate ODE system. For instance, we can easily prove that for Hamiltonian systems it does not preserve in general the symplectic structure of the phase space, as shown in Proposition 1.
Proposition 1.
The fixed point discretization is not symplectic for general Hamiltonian systems.
Proof.
We consider an Hamiltonian system defined by the Hamiltonian $H=H(q^{(j)},p^{(j)})$ for $j=1,\dots,m$ and the canonical coordinates $\overline{x}=(\overline{q},\overline{p})\in\mathbb{R}^{2m}$. The Hamilton equations are
[…]
therefore the fixed point discretization is given by
[…]
It is sufficient whether the symplecticity condition is satisfied, that is
$J^{T}MJ=M\,,\quad M=\begin{pmatrix}0&\mathbb{1}_{m\times m}\\ -\mathbb{1}_{m\times m}&0\end{pmatrix}\,$, (3)
where J is the Jacobian of the system (2). A simple calculation is sufficient to see that the symplecticity condition holds only for the specific class of Hamiltonian that satisfy
$C^{T}A=(C^{T}A)^{T}\,,\quad A^{T}B=(A^{T}B)^{T}\,,\quad C^{T}B-A^{T}A=\mathbb{1}_{m\times m}\,$, (4)
where
[…]
As an example, for one-dimensional Hamiltonian $H=H(q,p)$, $q,p\in\mathbb{R}$, the fixed point discretization is symplectic only for Hamiltonian with unitary Hessian.
Nevertheless, our aim is not to obtain a discrete approximation of the ODEs system, but rather to define an alternative discrete version associated to the continuous system and to explore connections between regular and chaotic behavior of both systems. The discrete system can share some properties with the continuous one (e.g. fixed points); on the other hand, it can exhibit a richer dynamics. Furthermore, the discretization we define can be thought as a novel chaotification procedure that associates a Li-Yorke chaotic discrete system to an integrable continuous one, under suitable constraints on the parameters.
Goal of this paper is to show that a large class of continuous dynamical systems has a fixed point discretization that is Li-Yorke chaotic for non trivial choices on the parameters, regardless the integrability or chaoticity of the continuous system. The starting point will be the discrete map defined in Disca and Coscia [2024, 2025]; in this work we relax the original constraints on the coefficients and define the Ziegler discrete map as
@@
The map (6) is obtained by discretizing in the above sense the equations of motion of the Ziegler pendulum in the sense of Polekhin Polekhin [2024] for a choice on the parameters leading to non-Hamiltonian integrability. From a physical point of view, parameters $\alpha$ and $\sigma$ involve masses and lengths of the pendulum, $\beta$ models an angular elastic potential and $\gamma$ represents the modulus of an external non-conservative force acting on the system. In the following we ignore this physical interpretation, focusing on general ranges on the parameters and emerging of chaotic behavior.
We also define the modular Ziegler discrete map as
@@ 3 Topological transitivity for the Ziegler map
We recall the definition of chaos in the sense of Devaney Devaney [2021], that is formulated through topological transitivity and density of periodic points.
@@
It is worth to notice that the original definition of Devaney was also considering sensitive dependence on initial data, a typical signature of chaotic systems; this condition has been proved to be redundant in Banks et al. [1992].
In this section we reintroduce the Ziegler map as defined in Disca and Coscia [2024, 2025], that is
@@
In Disca and Coscia [2025] it is proved that the map (8) is not Devaney chaotic for $\alpha<0$, $\tilde{\alpha}>0$, $\beta>0$, $\gamma\neq 0$, since it has not dense sets of periodic points; in particular, the map has not even periodic points and the sets of odd periodic points are not dense in $\mathbb{R}^{4}$. We show that the remaining necessary condition for the Devaney chaos is not satisfied for a symmetric choice on the parameters.
@@ 4 Chaotic discretization theorems
and derive our results for the map (30) with arbitrary small values of $\varepsilon_{1}$, $\varepsilon_{2}$; naturally, (30) reduces to (6) for $\varepsilon_{1},\varepsilon_{2}\to 0^{+}$. From a physical point of view, the coefficients $\varepsilon_{1}$, $\varepsilon_{2}$ model a (small) negative-dumping mechanism acting on the system.
We divide the proof in a few steps, by analyzing separately the sufficient conditions stated in Theorem 2 and deriving suitable constraints on the parameters.
@@
Fixed points of (6) are known from Disca and Coscia [2024]; by considering the new definitions given in (6) and (7), we can write them through the solutions of
@@
Let $J$ be the Jacobian of (30). $\forall\varepsilon_{1},\varepsilon_{2}>0$, there exist $\overline{\sigma}(\varepsilon_{1},\varepsilon_{2})\in\mathbb{R}$ and a fixed point $p_{0}$ such that the matrix $J^{T}J(p_{0})$ has all eigenvalues greater than $1$, provided that $|\beta|>\frac{1}{\sqrt{2}}$ and $|\sigma|>\overline{\sigma}$.
@@
We have $\lambda_{\pm}>1$ for a sufficiently great value for the modulus of $\sigma$; furthermore, $\lambda_{\beta}>1$ if $|\beta|>\frac{1}{\sqrt{2}}$. Finally, $\lambda_{1}>1$ $\forall\varepsilon_{1}>0$. We conclude that, $\forall\varepsilon_{1,2}>0$ and $|\beta|>\frac{1}{\sqrt{2}}$, there exists a threshold parameter $\overline{\sigma}(\varepsilon_{1},\varepsilon_{2})\in\mathbb{R}$ such that $\lambda_{1},\lambda_{\beta},\lambda_{\pm}>1$, provided that $|\sigma|>\overline{\sigma}$. ∎
@@
Given Lemma 3, we can avoid the need to construct a proper neighborhood of $p_{0}$. Furthermore, the map (6) has not even periodic points Disca and Coscia [2025], so we can show the existence of a snap-back for the point $p_{0}$ by proving the existence of solutions for the equation $f^{(2)}(q)=p$; since it cannot exists a periodic point of period 2, it follows that $q_{0}\neq p_{0}$.
@@
Then, for $\varepsilon_{1,2}$ sufficiently small, $\exists\overline{\sigma}(\varepsilon_{1},\varepsilon_{2})\in\mathbb{R}$ such that (43) is chaotic in the sense of Li-Yorke $\forall\sigma$ such that $|\sigma|>\overline{\sigma}$.
@@
where $\varepsilon_{1,2}>0$, $\alpha>0$, $\beta>0$, $\gamma\in(-2\pi,0)$, $\sigma<-1$ and
@@
Then, for $\varepsilon_{1,2}$ sufficiently small, $\exists\overline{\sigma}(\varepsilon_{1},\varepsilon_{2})\in\mathbb{R}$ such that (45) is chaotic in the sense of Li-Yorke $\forall\sigma$ such that $|\sigma|>\overline{\sigma}$.
@@
and $B^{(j)}$ has at least two distinct real roots. Then, for $\varepsilon_{1,2}$ sufficiently small, $\exists\overline{\alpha}(\varepsilon_{1},\varepsilon_{2})\in\mathbb{R}$ such that (47) is chaotic in the sense of Li-Yorke $\forall\alpha$ such that $|\alpha|>\overline{\alpha}$.
@@
Theorem 5 is particularly relevant in the context of nonlinear oscillators Kovačić [2020], when they are perturbed by periodic external forces, defined by polynomial potentials and are subject to linear velocity-dependent perturbations.
As an explicit example of application, Theorem 3 - 5 can be applied to proper functional extensions of discrete models for cardiac arrhythmia, such as the Guevara map proposed in Guevara and Glass [1982].
4.1 Qualitative remarks on scrambled sets and perturbations
In this section we outline a heuristic argument regarding potential formal connections between Li-Yorke chaos and KAM theory for quasi-integrable dynamical systems. The following discussion is meant to be a roadmap for future research works in the field.
@@ 4.1 Qualitative remarks on scrambled sets and perturbations
In answering this question, a curious connection with the KAM theorem Kolmogorov [1979], Arnold [1963], Moser [1966] can be conjectured. From KAM theory, we know that a quasi-integrable system exhibits invariant tori in the phase space that are not destroyed by the presence of a perturbation; when acting on the system, the perturbation destroys some tori, while other ones survive but are deformed, according to specific non-resonance conditions. In the context of chaos theory, it might be argued that a perturbation acts on the scrambled set of a Li-Yorke chaotic map in a similar way, in the sense that they could exist specific sufficient conditions such that the scrambled set of a Li-Yorke map survives under a small perturbation. If this is true, it would be possible to deduce chaos for a discrete system in general through perturbative arguments.
@@ 5.1 Looking for a strange attractor
In Figure 3 - 4 we set $\tilde{\alpha}\neq\alpha$. It is recognized the arising of a strange attractor for the system, that exhibits a fractal structure analogous to the Hénon attractor Hénon [1976]. This attractor is confined to the projection planes $(x,z)$ and $(y,\omega)$, while it is replaced by completely random motion by taking projections on the planes $(x,y)$ and $(z,\omega)$, as shown in Figure 5.
Figure 3: Confined strange attractor for the projection onto the plane $(x,z)$, non-symmetric case $\tilde{\alpha}\neq-\alpha$. The simulation runs for $10^{6}$ iterations. Parameters: $\alpha=-2.5$, $\tilde{\alpha}=2$, $\beta=0.8$, $\gamma=10$. Initial conditions: $x_{0}=0.2$, $y_{0}=0.1$, $z_{0}=0.8$, $\omega_{0}=0.5$. a) Full attractor. b) Magnification of a) related to the red rectangle. c) Magnification of a) related to the green rectangle.
Figure 4: Confined strange attractor for the projection onto the plane $(y,\omega)$, non-symmetric case $\tilde{\alpha}\neq-\alpha$. The simulation runs for $10^{6}$ iterations. Parameters: $\alpha=-2.5$, $\tilde{\alpha}=2$, $\beta=0.8$, $\gamma=10$. Initial conditions: $x_{0}=0.2$, $y_{0}=0.1$, $z_{0}=0.8$, $\omega_{0}=0.5$. a) Full attractor. b) Magnification of a) related to the red rectangle. c) Magnification of a) related to the green rectangle.
@@
Now, we briefly focus on the case $\gamma=0$. By recalling the physical origin of the system (6), the term $\gamma$ represents an external non-conservative force acting on an Hamiltonian system; by setting $\gamma=0$, the map comes out from an integrable continuous system whose phase space is foliated in invariant tori, due to Liouville-Arnold theorem Liouville [1855], Arnold [1989]. Then, it is worth to see whether the strange attractor shown in Figure 3 - 4 survives for this unperturbed case.
@@
Figure 6: Confined strange attractor for the projection onto the plane $(x,z)$, non-symmetric case $\tilde{\alpha}\neq-\alpha$. The simulation runs for $10^{6}$ iterations. Parameters: $\alpha=-2.5$, $\tilde{\alpha}=2$, $\beta=0.8$, $\gamma=0$. Initial conditions: $x_{0}=0.2$, $y_{0}=0.1$, $z_{0}=0.8$, $\omega_{0}=0.5$. a) Full attractor. b) Magnification of a) related to the red rectangle.
Figure 7: Confined strange attractor for the projection onto the plane $(y,\omega)$, non-symmetric case $\tilde{\alpha}\neq-\alpha$. The simulation runs for $10^{6}$ iterations. Parameters: $\alpha=-2.5$, $\tilde{\alpha}=2$, $\beta=0.8$, $\gamma=0$. Initial conditions: $x_{0}=0.2$, $y_{0}=0.1$, $z_{0}=0.8$, $\omega_{0}=0.5$. a) Full attractor. b) Magnification of a) related to the red rectangle.
In Figure 8 they are shown projections on the planes $(x,z)$ and $(y,\omega)$ of two trajectories with very close initial conditions, giving a further confirmation of the sensitive dependence on initial data for the system. The sensitive dependence of the system shown in Figure 8 will found another confirm in Section 5.2 through the computation of positive Lyapunov exponents.
Figure 8: Confined strange attractor for two very close initial conditions. Parameters: $\alpha=-3.5$, $\tilde{\alpha}=2$, $\beta=0.8$, $\gamma=-4$. Initial conditions, blue orbit: $x_{0}=1.4$, $y_{0}=0.8$, $z_{0}=2.7$, $\omega_{0}=2.34$. Initial conditions, green orbit: $x_{0}=1.4001$, $y_{0}=0.7999$, $z_{0}=2.7001$, $\omega_{0}=2.3399$. The simulation runs for $10^{5}$ iterations. a) Projection onto the plane $(x,z)$. b) Projection onto the plane $(y,\omega)$, magnification on the center.
5.2 Bifurcation diagrams and Lyapunov exponents
In this section we perform a bifurcation analysis of the system (7). We start with the following simple result.